# Could black hole thermodynamics play a role in black hole mergers?

### v1

@@ Abstract
Gravitational waves from binary black hole mergers yield values for both the black hole remnant mass $M$ and it’s spin $a$, with the $169$ $a$ values collected so far crowding significantly around their average $\bar{a}=0.6869\pm 0.087$. Could this crowding relate directly to the Davies phase transition point at $a=0.68125$ from black hole thermodynamics? I argue that a necessary challenge for such a connection requires a consistent application of the thermodynamic fluctuation theory that follows from black hole thermodynamics (BHT). Specifically, necessary are a correct choice of fluctuating variables, as well as thermal equilibrium between the event horizon at the Hawking temperature $\sim\mu K$ and the outside universe $\sim 3K$. I show that the former requirement follows in straightforward fashion from the BHT of the Kerr model, while the later requires an accretion disk following the Novikov-Thorne accretion disk model. I construct a thermodynamic fluctuation theory meeting both these requirements. My results open the possibility that black hole mergers are based on some dynamical model (not known to me) with a limiting attractor state at the Davies point.
Keywords: black hole mergers; black hole thermodynamics; thermodynamic fluctuation theory; phase transition; Davies singularity point; Novikov-Thorne accretion disk model;
@@ 1 Introduction
The detection of gravitational waves produced by the merger of binary stellar mass black holes (BH) [1, 2, 3, 4] has opened a new era in the study of BH’s. Present methods now allow for the measurement of both the mass $M$ and the spin angular momentum $J$ of the post-merger remnant BH’s. Following the lead of Wu et al. [5], I argue that black hole thermodynamics (BHT) [6, 7] might offer a framework for a striking pattern in the post-merger $(M,J)$ data. Measurements have found that remnant BH’s have dimensionless spin values
@@
close to the value of the Davies point (DP) $a=a^{*}=0.68125$ from black hole thermodynamics (BHT) [8]. The DP marks a point where fluctuations in $M$ diverge and thermodynamic stability changes from stable (at $a>a^{*}$) to unstable (at $a<a^{*}$). Is there actually some real causal physical connection between these observed post-merger spin data and the DP? So far as I am aware, this question has not yet been answered. Here I propose to make a good start based on thermodynamic fluctuation theory (TFT).
The DP is naturally grounded in the topic of TFT. To apply this theory here, two major challenges must be overcome. The first challenge is to determine which of the three conserved variables $(M,J,Q)$, with $Q$ the charge, actually fluctuate. A priori, it could be all three, or just one or two, with the other(s) drifting only slowly. A straightforward application of the Kerr BH model [9, 10] resolves this issue. I show that it suffices to consider only $\{M\}$ fluctuations. The second challenge is to establish thermal equilibrium between a stellar mass black hole, with it’s event horizon at the Hawking temperature $(\sim\mu K)$, and the rest of the universe at the cosmic background temperature ($\sim 3$ K). For this, I employ an intermediate co-rotating system to equalize the temperatures.
@@
Of course, the success that I report here applies only to the observed remnant spin data, because my TFT theory is based on that of just a single Kerr BH. But it is for this observed remnant spin data that we have what appears to be the interesting connection with the DP. Obviously, the full merger process calls for a dynamical theory certainly involving general relativity. But the case put forward here gives a possible context for a full dynamical theory: look for a theory that has the remnant end product of the merger process consisting of the DP as a limiting attractor state. This theory might involve the quantum mechanics that is an essential element of BHT.
@@ 2 The observational data
Look first at the observed remnant spin data motivating this paper. These data were taken in four sets of observing runs [1, 2, 3, 4], and are arranged in chronological order in Figure 1, with points colored brown, blue, green, and black, respectively. The observed post-merger spin data $a$ are dimensionless in the system of units given in [8]. On physical grounds, $a$ falls in the range $0\leq a<1$ [8], where $a=0$ corresponds to the non-spinning Schwarzschild black hole, and $a=1$ corresponds to the extremal limit where the black hole spins as fast as is physically possible.
@@
Figure 1(a) shows $a$ for the observed remnant spins reported so far. Visually, there is a clear tendency favoring spin values close to the DP, $a=a^{*}=0.68125$. The average observed value of $a$ is $\bar{a}=0.6869\pm 0.087$. There are $98/169$ data points above the DP, in the BHT stable regime, and $71/169$ data points below the DP, in the BHT unstable regime. $151/169$ of the data points ($90\%$) have their error bars encompassing $a^{*}$. There are 10/169 points with error bars completely above the DP, and 8/169 with error bars completely below the DP. These observed remnant spin data make a visual case that the DP might have some relevance to the observed data. For comparison, Figure 1(b) shows the corresponding remnant masses $m$, in solar units. The average value of $m$ is $\bar{m}=57.9\pm 32M_{\odot}$. No particular visual pattern in the distribution of $m$ values is evident.
@@
The Appendix adds some additional statistical observations.
@@ 3 The relation to thermodynamic stability
with $\hbar$ Planck’s constant divided by $2\pi$. The capitalized symbols $G$ and the solar mass $M_{\odot}$ are in MKS units. To be below the extremal value of rotation, $J$ must lie in the range $(-M^{2},M^{2})$. As emphasized by Bekenstein [14], a theory containing both $\hbar$ and $G$ has a quantum gravity flavor.
@@ 4 Basic thermodynamic fluctuation theory
Thermodynamic stability is usually presented in statistical mechanics textbooks [15, 16] in the context of thermodynamic fluctuation theory (TFT). TFT requires thermodynamic stability for its application. But, a priori this certainly does not force BH’s to live only in thermodynamically stable states. One can easily imagine that a BH could be held together by its powerful gravity regardless of its BHT stability. In this section, I present the basics of TFT, with new elements specific to this discussion added in subsequent sections.
@@
Here, $X^{U}$, $X^{A}$, and $X^{B}$ are the lists of values of the conserved fluctuating parameters of the systems in play here, including the whole universe $U$. The entropies of $U$, AA, and $B$ are denoted by $S^{U}(X^{A},X^{B})$, $S^{A}(X^{A})$, and $S^{B}(X^{B})$, respectively. The conservation laws yield $X^{U}=X^{A}+X^{B}$, which leads to
@@
Assuming thermodynamic stability, the equilibrium state has maximum total entropy = $S_{0}$, and coordinate values $X^{A}=X^{A}_{0}$. Expanding $S^{U}(X^{A})$ to second-order about $X^{A}_{0}$ in powers of $\Delta X^{A}=(X^{A}-X^{A}_{0})$, yields
@@ 6 A co-rotating environment is required for thermal equilibrium
A key finding of NT is that as $r\downarrow r_{isco}$, the internal stresses, the emitted heat, and the temperature are each predicted to go to zero; see Equation (11) of Reynolds [20]. Inside the ISCO, orbits cease to be stable and are predicted to plunge rapidly to the event horizon because the centrifugal force is no longer strong enough to resist gravity. There seems to be no consensus about the temperature between the ISCO and the event horizon, and this region is not physically observable. Nevertheless, there certainly seems a possibility that the temperature there could be low enough to equilibrate with the Hawking radiation. At present, we can only assume this, and we must, since else we can have no thermodynamic equilibrium, and TFT between system $A$ and $B$ fails, leaving us with no connection between the DP and the observational merger data other than coincidence.
@@ 9 Conclusion and future directions
In conclusion, gravitational wave observations have measured the spins $a$ of newly formed remnant stellar mass black holes (BH) resulting from the mergers of binary black hole systems. These spin values are found to cluster around the Davies singularity point (DP), which has $a=a^{*}=0.68125$. This suggests a significant connection between the results from gravity wave measurements and black hole thermodynamics (BHT).
@@
This paper emphasizes this numerical connection, including the latest (fourth) observed remnant spin data set. I have also contrasted the observed remnant spin data with that of the spin data from much older supermassive black holes. The later tend to spin very fast, with their BHT states overwhelmingly in the stable regime above the DP, and no clustering of spin values about any particular number.
@@
Most significantly, I have argued that the DP has its physical roots in thermodynamic fluctuation theory (TFT). In this context, the challenge is to construct a picture with the correct choice of fluctuating variables, as well as thermal equilibrium at the event horizons. I have met this challenge by arguing that in the Kerr model scenario we need consider only $\{M\}$ fluctuations, and proposing that thermal equilibrium at the event horizon could be achieved with an accretion disk following the Novikov-Thorne disk model (NT). I emphasize that if my approach fails, and is not replaced by a better one, we will be left with no connection between the data clustering in evidence here and BHT. In this case, the correlation shown in Fig. 1(a) will be mere coincidence, a disappointing outcome.
Of course, just establishing the relevance of the DP in this scenario still leaves us short of an actual dynamical mechanism by which the post-merger remnants form. General relativity seems to be the place to start looking for such dynamics. But the case put forward here suggests that it might be productive to try to build some dynamical theory that contains the DP as an attractor state for the final merged remnant black hole. This adds quantum mechanics to the theoretical mix with general relativity as a try to understanding BH mergers.
Finally, I list some future research projects suggested by this paper:
First, to continue to collect BH merger remnant data and see how $sr$ in Eq. (2) develops as the data collection builds.
Second, to find some microscopic justification for the BHT’s findings about the fluctuating variables: the domination of the $M$ fluctuations, the tiny $Q$ fluctuations, and the absence of a DP in the $J$ fluctuations. In these fluctuations, what kind of particles are involved? Does dark matter play a role?
Third, a further consideration of thermal equilibrium at the event horizon, an essential theme in my approach. Can the environment just outside the event horizon really get cold enough to match the Hawking temperature?
Finally, have any astronomical events been seen that might match a BH exploding as it spins down through the DP? Could a BH even get close enough to the DP for fluctuations in $M$ to get large enough to generate sufficient fluctuations? Or are the dominant fluctuating particles too granular to make this likely?
10 Appendix
In this Appendix, I present some additional statistical details that further a case that the observed remnant spin data $a$ cluster tightly around the DP. Let $N_{>}(a)$ be the number of observed remnant spin values above some general $a$. $N_{>}(a)$ is plotted against $a$ in Figure 6. In the data tables that I consulted [1, 2, 3, 4], $a$ is expressed as two digit numbers $0.00$, $0.01$, $0.02$, $\ldots 0.99$, $1.00$, and this gives my figure a somewhat down-stepping look. The red cross near the middle of Figure 6(b) has its horizontal bar set to half the number of particles, and its vertical bar set to $a^{*}$. Notice that the cross center falls close to the data curve. But the uncertainty in $\bar{a}$ is large enough that the cross center could, within error, fall on the data curve itself.
The following possible scenario is then consistent with the present observed remnant spin data: the BH remnants emerge from the mergers with spins very near $a^{*}$, much closer than the observed error bars in $a$. These error bars prevent us from directly seeing such a tight clustering. Regardless the issue of whether the BH’s are in the stable regime or not, we might expect that the observed values of $a$ are roughly equally split between above and below the DP, and with more data points a skew ratio near $sr=1/2$ would become more pronounced.
Refer to caption Refer to caption
Figure 6: $N_{>}$ plotted as a function of $a$, where $N_{>}$ is the number of observed remnant spin data points with spins larger than $a$: a) has all $169$ points included, and b) is focused on the DP. The red cross is centered at $a^{*}=0.68125$ and $N_{>}=169/2$.
Finally, I evaluate $\chi^{2}$, a standard measure of the quality of a fit between theory and data:
$\chi^{2}=\frac{1}{n}\sum_{i=1}^{n}\left[\frac{(a_{t}-a_{i})}{\sigma_{i}}\right]^{2},$ (25)
where $n$ is the number of data points, $a_{i}$ is the value of $a$ for the $i$’th data point, $\sigma_{i}$ is the length of the $i$’th error bar, and $a_{t}$ is the theoretical value for $a$ being tested. The values $a_{i}$ generally have different upper and lower error bars. To select between them, I set $\sigma_{i}$ to the upper error bar if $a_{i}$ is below $a_{t}$, and I set $\sigma_{i}$ to the lower error bar if $a_{i}$ is above $a_{t}$. The best fitting value of $a_{t}$ to the data is the one that minimizes $\chi^{2}$.
The full data set in play in Fig. 1(a) is a bit problematic for $\chi^{2}$ analysis. There are a number of outlier points having relatively small error bars, and deviating widely from the pack of points clustered around the DP. But for $a_{t}$ near $a^{*}$ (where $\chi^{2}$ turns out to be smallest) these outlier points contribute unreasonably large amounts to the sum, and run the danger of skewing the analysis unreasonably. Hence, for the $\chi^{2}$ analysis, I omitted the $18$ points whose error bars do not encompass $a^{*}$, leaving $151$ points in the data set. Figure 7 shows the resulting $\chi^{2}$ calculated from this pruned data set. This $\chi^{2}$ is minimized at $a_{t}=0.689$, very close to the Davies point $a^{*}$.
Figure 7: $\chi^{2}$ as a function of the theoretical value $a_{t}$ calculated from a pruned data set of $151$ observed remnant spin data points. The $a_{t}$ value that minimizes $\chi^{2}$ is $a_{t}=0.689$, indicated with the red arrow, and very near $a^{*}=0.68125$.

### v2

@@ Abstract
Gravitational waves detected from binary black hole mergers by the LIGO/Virgo/KAGRA collaboration yield values for both the black hole remnant mass $M$ and it’s spin $a$, with the $169$ $a$ values collected so far crowding significantly around their average $\bar{a}=0.6869\pm 0.0135$. Could this crowding relate directly to the Davies phase transition point at $a=0.68125$ from black hole thermodynamics? I argue that a necessary challenge for such a connection to obtain requires a consistent application of the thermodynamic fluctuation theory that follows from black hole thermodynamics (BHT). Specifically, necessary are a correct choice of fluctuating variables, as well as thermal equilibrium between the event horizon at the Hawking temperature $\sim\mu K$ and the outside universe $\sim 3K$. I show that the former requirement follows in straightforward fashion from the BHT of the Kerr model, while the later requires an accretion disk following the Novikov-Thorne accretion disk model. I construct a thermodynamic fluctuation theory meeting both these requirements. My results open the possibility that black hole mergers are based on some dynamical model (not known to me) with a limiting attractor state at the Davies point.
Keywords: black hole mergers; black hole thermodynamics; thermodynamic fluctuation theory; phase transition; Davies singularity point; Novikov-Thorne accretion disk model; LIGO; Virgo; KAGRA;
@@ 1 Introduction
The detection of gravitational waves produced by the merger of binary stellar mass black holes (BH) by the LIGO/Virgo/KAGRA collaboration [1, 2, 3, 4] has opened a new era in the study of BH’s. Present methods now allow for the measurement of both the mass $M$ and the spin angular momentum $J$ of the post-merger remnant BH’s. Following the lead of Wu et al. [5], I argue that black hole thermodynamics (BHT) [6, 7] might offer a framework for a striking pattern in the post-merger $(M,J)$ data. Measurements have found that remnant BH’s have dimensionless spin values
@@
close to the value of the Davies point (DP) $a=a^{*}=0.68125$ from black hole thermodynamics (BHT) [8]. The DP marks a point where fluctuations in $M$ diverge and thermodynamic stability changes from stable (at $a>a^{*}$) to unstable (at $a<a^{*}$). Is there actually some real causal physical connection between these observed post-merger spin data and the DP? So far as I am aware, this question has not yet been answered definitively. Here I propose to make a good start based on BHT, in the form of its thermodynamic fluctuation theory (TFT).
To determine the $(M,J)$ values of a remnant BH, the LIGO/Virgo/KAGRA team first measure the gravitational wave “chirp” emitted by the source merger event. Measured and theoretically computed chirps are then compared to determine the best match between them. The theoretically computed chirps result from numerical relativity calculations, tabulated in a preprepared catalog. These catalog entries include the computed remnant values $(M,J)$ for each chirp. Numerical relativity consists of numerical solutions of the Einstein field equations, subject to initial conditions. In practice, this powerful solution method works beautifully. However, it offers only numbers, with no general principle for why the bunching of the $a$ values at the DP occurs. My hope is that bringing in BHT and its phase transitions can lead to an eventual solution.
To apply TFT in the BH scenario, two major challenges must be overcome. The first challenge is to determine which of the three conserved variables $(M,J,Q)$, with $Q$ the charge, actually fluctuate. A priori, it could be all three, or just one or two, with the other(s) drifting only slowly. A straightforward application of the Kerr BH model [9, 10] resolves this issue. I show that it suffices to consider only $\{M\}$ fluctuations. The second challenge is to establish thermal equilibrium between a stellar mass black hole, with it’s event horizon at the Hawking temperature $(\sim\mu K)$, and the rest of the universe at the cosmic background temperature ($\sim 3$ K). For this, I employ an intermediate co-rotating system to equalize the temperatures.
@@
Obviously, the full merger process calls for a full dynamical theory. The case put forward here gives a possible context for such a theory: look for a theory that has the remnant end product of the merger process consisting of the DP as a limiting attractor state.
@@ 2 The observational data
Look first at the observed remnant spin data motivating this paper. These data were taken in four sets of observing runs [1, 2, 3, 4], and are arranged in rough chronological order in Figure 1, with points colored brown, blue, green, and black, respectively. The observed post-merger spin data $a$ are dimensionless in the system of units given in [8]. On physical grounds, $a$ falls in the range $0\leq a<1$ [8], where $a=0$ corresponds to the non-spinning Schwarzschild black hole, and $a=1$ corresponds to the extremal limit where the black hole spins as fast as is physically possible.
@@
Figure 1(a) shows $a$ for the observed remnant spins reported so far. Visually, there is a clear tendency favoring spin values close to the DP, $a=a^{*}=0.68125$. The average observed value of $a$ is $\bar{a}=0.6869\pm 0.0135$, where $0.0135$ is the margin of error yielding a 95% confidence interval. The DP value of $0.68125$ falls into this observed range, so it is consistent with the measured remnant values.
There are $98/169$ data points above the DP, in the BHT stable regime, and $71/169$ data points below the DP, in the BHT unstable regime. $151/169$ of the data points ($90\%$) have their error bars encompassing $a^{*}$. There are 10/169 points with error bars completely above the DP, and 8/169 with error bars completely below the DP. These observed remnant spin data make a visual case that the DP might have some relevance to the observed data.
For comparison, Figure 1(b) shows the corresponding remnant masses $m$, in solar units $M_{\odot}$. The average value of $m$ is $\bar{m}=57.9\pm 32M_{\odot}$. No particular visual pattern in the distribution of $m$ values is evident.
@@ 3 The relation to thermodynamic stability
with $\hbar$ Planck’s constant divided by $2\pi$. The capitalized symbols $G$ and the solar mass $M_{\odot}$ are in MKS units. To be slower than the extremal value of rotation, $J$ must lie in the range $(-M^{2},M^{2})$. As emphasized by Bekenstein [14], a theory containing both $\hbar$ and $G$ has a quantum gravity flavor.
@@ 4 Basic thermodynamic fluctuation theory
Thermodynamic stability is usually presented in statistical mechanics textbooks [15, 16] in the context of thermodynamic fluctuation theory (TFT). TFT requires thermodynamic stability for its application. But, a priori, this certainly does not force BH’s to live only in thermodynamically stable states. One can easily imagine that a BH could be held together by its powerful gravity regardless of its BHT stability. In this section, I present the basics of TFT, with new elements specific to this discussion added in subsequent sections.
@@
Here, $X^{U}$, $X^{A}$, and $X^{B}$ are the lists of values of the conserved fluctuating parameters of the systems in play here. The entropies of $U$, AA, and $B$ are denoted by $S^{U}(X^{A},X^{B})$, $S^{A}(X^{A})$, and $S^{B}(X^{B})$, respectively. The conservation laws yield $X^{U}=X^{A}+X^{B}$, which leads to
@@
Assuming thermodynamic stability, the equilibrium state has maximum total entropy = $S_{0}$, at coordinate values $X^{A}=X^{A}_{0}$. Expanding $S^{U}(X^{A})$ to second-order about $X^{A}_{0}$ in powers of $\Delta X^{A}=(X^{A}-X^{A}_{0})$, yields
@@ 6 A co-rotating environment is required for thermal equilibrium
A key finding of NT is that as $r\downarrow r_{isco}$, the internal stresses, the emitted heat, and the temperature are each predicted to go to zero; see Equation (11) of Reynolds [20]. Inside the ISCO, orbits cease to be stable and are predicted to plunge rapidly to the event horizon because the centrifugal force is no longer strong enough to resist gravity. There seems to be no consensus about the temperature between the ISCO and the event horizon, and this region is not physically observable. Nevertheless, the possibility does arise that the temperature there could be low enough to equilibrate with the Hawking radiation. At present, we can only assume this, and we must, since else we can have no thermodynamic equilibrium, and TFT between system $A$ and $B$ fails, leaving us with no connection between the DP and the observational merger data other than coincidence.
@@ 9 Conclusion
In conclusion, gravitational wave observations have measured the spins $a$ of newly formed remnant stellar mass black holes (BH) resulting from the mergers of binary black hole systems. These measured spin values are found to cluster around the Davies singularity point (DP), which has $a=a^{*}=0.68125$. This suggests a significant connection between the results from gravity wave measurements and black hole thermodynamics (BHT).
@@
This paper emphasizes this numerical connection, including the latest (fourth) observed remnant spin data set. I have also contrasted the observed remnant spin data with that of the spin data from much older supermassive black holes. The later tend to spin very fast, with their black hole thermodynamic (BHT) states overwhelmingly in the stable regime above the DP, and no clustering of spin values about any particular number.
@@
Most significantly, I have argued that the DP has its physical roots in thermodynamic fluctuation theory (TFT). In this context, the challenge is to construct a picture with the correct choice of fluctuating variables, as well as thermal equilibrium at the event horizons. I have met this challenge by arguing that in the Kerr model scenario we need consider only $\{M\}$ fluctuations, and by proposing that thermal equilibrium at the event horizon could be achieved with an accretion disk following the Novikov-Thorne disk model (NT). I emphasize that if my approach fails, and is not replaced by a better one, we will be left with no connection between the data clustering in evidence here and BHT. In this case, the correlation shown in Fig. 1(a) will be mere coincidence, a disappointing outcome.
Of course, just establishing the relevance of the DP in this scenario still leaves us short of an actual dynamical mechanism by which the post-merger remnants form. Numerical relativity certainly deals with such a mechanism, and its computations yield the measured remnant values of $a$ analyzed here. But it misses the connection to the DP, and the reason why the DP appears to be an attractor state for the final merged remnant black holes. I hope that the introduction of the thermodynamic fluctuations present in BHT can shed some insight into these processes.
10 Acknowledgements
I thank Andrey Skripnikov for help with the statistics. I also thank the late George Skestos for inspiring me to write about observational gravity with his support, and with his intense questions about gravity.